3. A 25.00 mL 0.0250 M calcium carbonate sample is titrated with 14.3 mM EDTA solution. Both sample and titrant are buffered at pH = 10.0 a. What is the titration reaction and what is the value of the conditional formation constant? b. What is the calcium ion concentration when 12.35 mL of titrant has been added? c. What is pCa2+ at the equivalence volume?
Answer #1

Given that Volume of Ca+2 solution = 25.00 mL Conc. of Cat2 solution = 0.0250 M Conc. of EDTA = 14.3 mm = 0.0143 M. logk = 10

(b) Addition of 12.35 mL of EDTA mmols of Ca+2 = M*V = 0.0250 M* 25.00 mL = 0.625 mmol mmols of EDTA = M*V = 0.0143 M* 12.35

(c) At Equivalence point: Ca+2 forms 1:1 complex with EDTA At equivalence point, Number of moles of Ca+2 = Number of moles of